A small ball is dropped from a balloon moving vertically up at a speed 12 m/s when the balloon is a height of 30.4 m. After what time the stone will reach the ground? g= 10 m/s²

Grade:9

1 Answers

Pete
5 years ago
As ball is with balloon initially when it is dropped so the velocity of ball was 12m/s in the upward direction now as we know that gravitational force acts downwards (taking up as +ve direction and down as -ve direction) so the ball will ultimately fall towards earth. So here as we know the initial speed of ball and value of gravitational acceleration along with initial height of release so we can use " h=ut+(at²)/2".:-30.4=12t-(gt²)/2. { g=10m/s² and in -ve direction also height is in downward}Putting the values we`ll get a quadratic in t and on solving it we get:t= 3.95 and -1.54 ..... As time cant be negative so "t=3.95 seconds===>ans."Thanks.Pete